OpenStudy (anonymous):

Claudia is on a motorcycle trip. She checked the odometer at the beginning of the trip, and it showed 16,789 miles. After 15 days it showed 22,189 miles. If Claudia rode the same number of miles each day, how many miles did she ride in a day?

OpenStudy (anonymous):

360* miles a day

OpenStudy (callisto):

total miles she rode = 22189 - 16789 = 5400 miles she rode a day = 5400/15 = 360 miles a day

OpenStudy (anonymous):

Medals for me and volley Callisto please, for helping you?(: